Timetable Monday, September 09

Time  Room LG 1, HS 1  Room LG 1, HS 2  Room LG 1, HS 304

08.00 – 09.00 Registration in Room LG 1, 110a

09.00 – 09.30 Opening Session in Room LG 1, HS 1

09.30 – 10.15 Plenary Talk (Chair: Diethard Pallaschke) Stefan Rolewicz: Abstract convex analysis in metric spaces

10.15 – 10.45  Coffee break

10.45 – 11.30 Plenary talk (Chair: Helmut Maurer) Richard Epenoy: Decompositioncoordination techniques in optimal control for computing multi-phases interplanetary trajectories of low-thrust probes

11.35 – 12.05 (Chair: Andreas Fischer) Jacek Gondzio: Preconditioning indefinite systems in interior point methods for optimization (Chair: Olga Kostyukova) Holger Irrgang: About gaps between optimal control problems and their Rubio generalizations (Chair: Arnd Rösch) Tim Voetmann: Weak regularization for contact problems (with friction)

12.05 – 12.35 Petra Huhn: Probabilistic analysis of interior-point methods Werner H. Schmidt: Existence of classical solutions or chattering controls for an optimal aeroassisted orbital transfer Alexander Kolpakov: On the discrete convex combinations problem

12.35 – 14.00  Lunch

14.00 - 14.45 Plenary talk (Chair: Rainer Tichatschke) Hubertus Th. Jongen: On minmax digraphs in global optimization

Timetable Monday, September 09 (continued)

Time  Room LG 1, HS 1  Room LG 1, HS 2  Room LG 1, HS 304

14.50 – 15.20 (Chair: Alexander Ioffe) Rolf Klötzler: On the geological structure of a riverbed (Chair: Marco Antonio López) Kristin Winkler: On multicriterial optimization with C(T)-valued functions (Chair: René Henrion) Andrzej Fryszkowski: On the solution set of the differential inclusions −Δ u ∈ F (t, u) in H_{0}^{1}(Ω)

15.20 – 15.50 Marcus Wagner: ε-optimality conditions for nonconvex multidimensional control problems Andreas Löhne: Scalarization functionals on set spaces Zdzislaw Naniewicz: Semicoercive variational-hemivariational inequalities with unilateral growth condition

15.50 – 16.15  Coffee break

16.15 – 16.45 (Chair: Vladimir F. Demyanov) Christian Grossmann: Path-follo-wing Newton methods and penalty-bar-rier results (Chair: Darinka Dentcheva) Elza Farkhi: Set-valued stability in differential inclusions with relaxed Lipschitz right-hand sides (Chair: Anatolij S. Antipin) Marc Steinbach: Custom sparse solvers for multistage stochastic programs

16.45 – 17.15 Moritz Diehl: A Newton-type method for real-time optimization in nonlinear model predictive control Ewa M. Bednarczuk: Order-Lipschit-zian properties of multifunctions with applications to stability of minimal points Mirjam Dür: Probabilistic subset selection in branch-and-bound algorithms

17.15 – 17.45 Marek Smietanski: New versions of approximate Newton methods for nonsmooth equations Jiri Outrata: On calmness of a class of multifunctions Housnaa Zidani: Numerical schemes for the stochastic HJB equation

19.30  Reception